OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)?

OpenStudy (anonymous):

anwsers_ x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ -4}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)